OpenStudy (anonymous):

Airplane A travels 1400 km at a certain speed. Plane B travels 1000 km at a speed 50 km/h faster than plane A in 3 hrs less time. Find the speed of each plane.

hero (hero):

I used to hate these..now they're my favorite...

hero (hero):

Are you there?

OpenStudy (anonymous):

I am here. I usually can figure these out, but having trouble with this one.

hero (hero):

I see....sorry for taking so long...would you like me to post a classroom and show you?

hero (hero):

The only part that I saw was "Report Abuse". Did you copy this from somewhere else? lol

hero (hero):

No disrespect Duke, but I was going to show nettej94 a much easier way to solve it that would involve less confusion.

hero (hero):

Yes, I know Duke.

OpenStudy (anonymous):

let da = 1400 km let Va = speed of plane A let ta = time a let Vb = speed of plane B let db = 1000 km let tb = time b db = Vb*tb 1000 km = (Va + 50 km/h)(da/Va - 3h) 1000 km = (Va + 50 km/h)(1400km/Va - 3h) now solve for Va add 50 km/h to get Vb oh a relationship between da, Va and ta may be helpful da = Va*ta db = Vb*tb

OpenStudy (anonymous):

I did copy it from somewhere else.

OpenStudy (anonymous):

open study glitchy right now on my CPU. I post then go back and see nothing I did copy my work from a page that I posted the answer remember many ways to solve this problem if it was originally a physics problem

OpenStudy (anonymous):

going to log out and shut down maybe site will work for me tomorrow

hero (hero):

I used to have those kinds of problems with OS....not anymore...

OpenStudy (anonymous):

A: r t = d, r = d/t, r=1400/t B: (r + 50) (t - 3) =1000 Replace r in the above equation with 1400/t and solve for t. t = -12 and 7. Use t = 7 and solve r=1400/t for r. r = 200 A: r = 200 mph B: r + 50 = 250 mph Verify B:'s equation: (200 + 50)(7 - 3) = 1000 ? 1000 = 1000 yes